OpenStudy (anonymous):

After a severe storm, three sisters, April, May and June stood on their front porch and noticed that the tree in their front yard was leaning 3 degrees from vertical toward the house. From the porch, which is 101 feet away from the base of the tree, they noticed that the angle of elevation to the top of the tree was 32 degrees. What is the height of the tree?.

OpenStudy (anonymous):

We have: AT cos 32 = AB CT cos (90-3) = CT cos 87 = BC Since all the angles in a triangle add up to 180 deg, angle ATC = 180 - angle TAC - angle ACT = 180 - 32- 87 = 61 deg Now, by the sine rule, sin ATC / AC = sin TAC / CT which gives CT = AC*sin TAC / sin ATC or CT = 101*sin 32 / sin 61 or CT = 101*0.53/0.875 = 61.194 ft The tree is 61.194 ft high